OpenStudy (anonymous):

How would I solve this problem by completing the square? 2y squared + 2y = -1

jimthompson5910 (jim_thompson5910):

2y^2 + 2y = -1 2(y^2 + y) = -1 2(y^2 + y + 1/4 - 1/4) = -1 2( (y^2 + y + 1/4) - 1/4) = -1 2( (y + 1/2)^2 - 1/4) = -1 2(y + 1/2)^2 - 2(1/4) = -1 2(y + 1/2)^2 - 2/4 = -1 2(y + 1/2)^2 - 1/2 = -1 I'll let you take over from here

OpenStudy (anonymous):

so far I have (2y + 1\4) - 1\2 = -1

OpenStudy (anonymous):

would I add the 1 to the left hand side so a 0 would be on the right side and then do the quadratic formula?

jimthompson5910 (jim_thompson5910):

2(y + 1/2)^2 - 1/2 = -1 2(y + 1/2)^2 = -1+1/2 2(y + 1/2)^2 = -1/2 (y + 1/2)^2 = (-1/2)*(1/2) (y + 1/2)^2 = -1/4

OpenStudy (anonymous):

so then y = -1\2?

jimthompson5910 (jim_thompson5910):

(y + 1/2)^2 = -1/4 y + 1/2 = sqrt(-1/4) or y + 1/2 = -sqrt(-1/4) y + 1/2 = (1/2)*i or y + 1/2 = -(1/2)*i .... .... y = ??? or y = ???

OpenStudy (anonymous):

so then that makes y = 1\2, -1\2, 1\2i, or -1\2i

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

but isn't the square root of 1\4 1\2?

jimthompson5910 (jim_thompson5910):

y + 1/2 = (1/2)*i or y + 1/2 = -(1/2)*i y = - 1/2 + (1/2)*i or y = - 1/2 - (1/2)*i

jimthompson5910 (jim_thompson5910):

sqrt(-1/4) = (1/2)*i